# Deep Unfolding with Approximated Computations for Rapid Optimization

### v1

@@ I Introduction
Iterative optimizers lie at the heart of numerous signal processing and communication tasks, including detection, estimation, and control [2]. These methods provide principled solutions to tasks that can be mathematically formulated as optimization problems. However, various applications, ranging from wireless communications [3] to real-time image processing [4] and edge computing [5], impose tight latency constraints that can render conventional iterative solvers impractical. In such settings, the time available to compute a solution is severely limited, and the use of classical optimization algorithms, which typically require tens to hundreds of iterations to converge, becomes a bottleneck. This pressing need for rapid and efficient solvers motivates the development of alternative approaches that can deliver high-quality solutions within tight computational budgets.
@@
Traditionally, the performance of iterative optimization algorithms has been studied from a convergence perspective, focusing on their asymptotic behavior as the number of iterations grows [6]. In this classical view, the choice of algorithmic hyperparameters, such as step sizes, penalty weights, or regularization strengths, is typically analyzed under assumptions that ensure theoretical convergence guarantees, often irrespective of their exact numerical values. However, in practice, these hyperparameters critically affect not only the convergence rate but also the computational efficiency of the solver. Techniques such as adaptive step size selection or backtracking line search are commonly employed to accelerate convergence in terms of iterations [7, Ch. 3], [8]. Yet, these methods often increase per-iteration complexity, as they require evaluating multiple candidate updates, ultimately leading to increased overall latency, which is the very constraint that practical systems aim to minimize. Other methods include derivative-free optimization, where surrogate models and adaptive restart mechanisms are used to enhance robustness and efficiency under expensive or noisy function evaluations [9]. However, these solvers typically require multiple sequential objective evaluations and do not address the per-iteration computational cost, as they rely on full-precision model evaluations.
@@
Recent years have witnessed a growing interest in leveraging deep learning tools to enhance optimization algorithms in various applications [10, 11, 12], giving rise to the paradigm of learning to optimize [13]. Among the most structured and interpretable approaches within this paradigm is deep unfolding [14], which systematically transforms iterative optimization algorithms into trainable neural network architectures. In deep unfolding, each iteration of the original solver is mapped to a layer in a feedforward network, with a common approach treating its algorithmic hyperparameters as learnable parameters [15]. A key advantage of this approach lies in its ability to optimize the optimizer itself under a fixed computational budget, allowing one to preserve the interpretability and domain-specific structure of the original method while tuning its behavior to match the statistics of the data [16]. This has been shown to enable reliable operation with a limited number of iterations of unfolded optimizers adopted in various domains, including wireless communications [17, 18], distributed optimization [19, 20], beamforming [21, 22, 23], integrated sensing and communications [24, 25], and outlier detection [26, 27], making the method highly suitable for latency-sensitive applications. Moreover, since the learning process is performed offline, no additional computational overhead is introduced at inference time, unlike adaptive schemes such as backtracking.
@@
In this paper, we propose a new class of learned optimization solvers based on deep unfolding, which compensates both for a limited number of iterations and for approximated computations within each iteration. Our key insight is that leveraging data allows designing unfolded architectures that effectively absorb and mitigate the performance loss introduced by deliberately simplified operations. This enables the design of optimization solvers that are both fast, due to a fixed and small number of unfolded iterations, and lightweight, due to replacing computationally expensive steps with efficient surrogates. Unlike standard deep unfolding methods that focus on reducing iteration count via, e.g., learned hyperparameters, our framework introduces a second axis of approximation: replacing costly operations within each iteration with low-complexity surrogates. These approximations are compensated for by expanding the learnable parameter space, allowing the optimizer to adaptively absorb their effects. This dual approximation (in both iteration depth and per-iteration complexity) enhances our approach’s applicability in latency- and resource-constrained environments. In doing so, we uncover an additional core benefit of deep unfolding: its capacity to tolerate and learn around structural approximations, thereby reducing both the number and cost of iterations without sacrificing solution quality.
@@
  • A new paradigm for learned approximated optimization: We introduce a principled framework for constructing unfolded approximated iterative optimizers, wherein selected steps of a classical iterative solver are replaced with lower-complexity operations. These approximate steps are embedded in a fixed-depth unfolded network whose parameters are optimized using training data, allowing the learned model to best match the data distribution while operating under strict latency constraints.
@@
  • Case study II — Robust principal component analysis: We further demonstrate the versatility of our approach by applying it to RPCA [29], a fundamental problem in signal processing. Using an unfolded version of the iterative thresholding algorithm with approximated singular value computations, we design a learned RPCA solver that maintains competitive recovery accuracy while dramatically reducing runtime.
@@ II-A Optimization Model
where $h_{{\bm{\theta}}^{\rm h}_{k}}:\mathcal{S}\times\mathcal{X}\mapsto\mathcal{S}$ denotes the update function applied at iteration $k$. This update is governed by solver-specific parameters ${\bm{\theta}}^{\rm h}_{k}$ (e.g., step sizes or momentum factors), referred to as hyperparameters, which guide the update steps but do not alter the optimization objective itself. While often fixed to constant values, which are either hand-tuned or derived via theoretical convergence criteria, hyperparameters play a critical role in determining the convergence speed, runtime and overall solution quality of the optimizer.
@@ III-A Rationale
As noted in C1 and C2, a major bottleneck in applying iterative optimization solvers to latency-sensitive applications stems from the variable and potentially large number of iterations required for convergence. To support rapid decision-making, it is thus essential to design optimization algorithms that operate with a fixed and relatively small number of iterations $K$, chosen in advance. A key enabler for such fixed-depth operation is the framework of deep unfolding, which has been shown to preserve the interpretable structure of classical solvers while learning iteration-specific hyperparameters from data [16]. In this paradigm, the iterative mapping (3) is viewed as a $K$-layer neural network, with each layer corresponding to a single iteration whose parameters are tuned offline for the target domain. By casting the hyperparameters $\{{\bm{\theta}}^{\rm h}_{k}\}_{k=1}^{K}$ as the trainable parameters of a discriminative machine learning model [30], one fully retains the interpretable operation of the original solver while achieving reliable performance within a small number of steps.
However, limiting the number of iterations alone does not fully resolve the latency bottleneck. As discussed in C3, the computational cost of each iteration may itself be prohibitive, particularly when the objective function involves costly gradient evaluations or matrix operations. To address this, we take a further step by replacing selected iterations with low-complexity approximated computations, designed to significantly reduce the per-iteration runtime. Of course, such approximations may degrade performance if left uncompensated. To maintain the reliability of the resulting optimization process, we extend the hyperparameter space of the unfolded solver in a way that does not increase computational burden (e.g., by using element-wise or structure-aware parameterizations), and we train the full approximated optimization procedure end-to-end using data. In this manner, we obtain an optimization model that is both interpretable and efficient, and can achieve strong performance with low latency and low computational overhead.
@@ III-B1 Increased Parameterization
This formulation increases the flexibility in learning iteration-specific dynamics, without increasing the number of floating-point operations, as the element-wise product is of the same computational order as scalar multiplication. This approach is reminiscent of variable metric algorithms with diagonal preconditioners [31]. In addition, it also preserves the semantics of descent-based optimization under mild regularity assumptions, as stated in the following proposition.
@@
where the approximation in $(a)$ holds for sufficiently small step sizes, and $(b)$ is obtained by substituting the update step from (5) and rearranging, with ${\rm diag}(\bm{\eta}_{k})$ being a diagonal matrix whose diagonal entries are $\bm{\eta}_{k}$. Since these entries are positive, then this diagonal matrix is positive-definite, and thus, $\mathcal{L}_{\rm o}({\bm{s}}^{(k+1)};{\bm{x}})\leq\mathcal{L}_{\rm o}({\bm{s}}^{(k)};{\bm{x}})$, i.e., (4) is a descent method. ∎
Proposition 1 demonstrates that increased abstractness through element-wise parameterization can be made to $(i)$ preserve computational complexity; $(ii)$ not compromise theoretical guarantees such as descent behavior; and $(iii)$ provide richer representational power to absorb approximations.
@@ III-B2 Constructing the Approximated Iterative Optimizer
Assume that the objective function $\mathcal{L}_{\rm o}({\bm{s}};{\bm{x}})$ is $\mu$-strongly convex and $L$-smooth with $L\geq\mu>0$. Let the step-size vector $\bm{\eta}_{k}$ satisfy $0\leq[\bm{\eta}_{k}]_{i}\leq\gamma<\frac{1}{L}$ for all $i$. In addition, suppose that for every $k\in\mathcal{K}^{\rm approx}$, the gradient approximation error satisfies
@@
for some sequence $\{\delta_{k}\}_{k=0}^{K-1}$. Then, for any minimizer $\bm{s}$$^{\star}$ of $\mathcal{L}_{\rm o}(\cdot;{\bm{x}})$, the error after $K$ steps satisfies:
$\|{\bm{s}}^{(K)}-{\bm{s}}^{\star}\|^{2}\leq(1-\gamma\mu)^{K}\|{\bm{s}}^{(0)}-{\bm{s}}^{\star}\|^{2}\hskip 28.45274pt$
$+\sum_{k\in\mathcal{K}^{\rm approx}}(1-\gamma\mu)^{K-k-1}\delta_{k}^{2}.$ (9)
@@
The first term in (2) corresponds to the standard convergence rate of gradient descent [32, Thm. 3.4]. The second term in (2) accumulates errors that are based on the deviations from the true gradient only over the selected iterations $\mathcal{K}^{\rm approx}$. This result shows that the final error scales with the magnitude of the skipped gradients and the approximation error $\delta_{k}$. The proposition indicates the potential usefulness of deliberately inducing approximated computations (for complexity reduction) with increased parameterization, since as long as the skipped gradients (balanced by the increased hyperparameters $\bm{\eta}_{k}$) are small, the error remains bounded and the method remains effective.
@@ III-B3 Training via Empirical Risk Minimization
To learn the parameters $\Theta\triangleq\{\Theta_{k}\}_{k=1}^{K}$ of the approximated iterative optimizer, we assume access to a representative dataset $\mathcal{D}$. The form of the training procedure depends on the nature of the available supervision.
@@
where $\mathcal{L}_{\rm task}(\cdot,\cdot)$ is a supervised loss function, such as mean squared error or cross-entropy, chosen according to the target application. The usage of supervised learning is most useful in settings where the optimization objective $\mathcal{L}_{\rm o}$ is a mathematically formulated proxy of the actual system task.
In both cases, the unfolded approximated optimizer is trained offline using standard gradient-based methods to minimize the empirical risk over $\Theta$. This allows the model to compensate for the approximations in both iteration count and complexity, yielding an efficient and reliable learned decision rule. The formulation of training an approximated unfolded optimizer via conventional deep learning methods, i.e., mini-batch sgd (sgd), is stated in Algorithm 1.
@@ III-C Discussion
The proposed methodology introduces a novel and flexible approach to accelerate iterative optimization by combining deep unfolding with approximated computations. While conventional uses of deep unfolding aim to reduce the number of iterations by learning solver hyperparameters from data, our framework extends this idea by deliberately introducing low-complexity approximations within the iterations themselves. These approximations are typically regarded as limitations that degrade performance; however, we show that when integrated into an unfolded architecture and paired with appropriate parameterization, the resulting model can be trained to absorb and compensate for the induced mismatches. In doing so, our methodology directly addresses all three challenges C1-C3. This dual approximation on both iteration depth and per-iteration complexity, is made tractable and effective through the trainability offered by deep unfolding. The increased abstractness in adding hyperparameters is designed not to come at the cost of increased complexity. Moreover, different abstractions, e.g., the usage of a per-parameter step-size rather than scalar ones as we do in our case studies, in fact preserve the property of descent methods as in (2[33].
Since our formulation is intentionally generic, many of the practical considerations and benefits of the proposed framework are revealed only when applied to specific use cases. In particular, while the goal is to reduce runtime complexity, the sources of computational cost, and thus the most effective approximation strategies, vary significantly across different iterative solvers. The choice of which computations to approximate and in which iterations to do so should generally be guided by predefined latency or hardware constraints, and tailored to the structure and demands of the application at hand. In the following sections, we illustrate how our general methodology can be instantiated in two representative signal processing tasks. These case studies demonstrate how approximations can be selected, quantify their computational impact, and empirically validate the ability of the learned optimizer to preserve or even improve performance.

### v2

@@ I Introduction
Iterative optimizers lie at the heart of numerous signal processing and communication tasks [2]. These methods provide principled solutions to tasks that can be mathematically formulated as optimization problems. However, various applications, ranging from wireless communications [3] to real-time image processing [4] and edge computing [5], impose tight latency constraints that can render conventional iterative solvers impractical. In such settings, the time available to compute a solution is severely limited, and the use of classical optimization algorithms, which typically require tens to hundreds of iterations to converge, becomes a bottleneck. This pressing need for rapid and efficient solvers motivates the development of alternative approaches that can deliver high-quality solutions within tight computational budgets.
@@
Traditionally, the performance of iterative optimization algorithms has been studied from a convergence perspective, focusing on their asymptotic behavior as the number of iterations grows [6]. In this classical view, the choice of algorithmic hyperparameters, such as step-sizes, penalty weights, or regularization strengths, is typically analyzed under assumptions that ensure theoretical convergence guarantees, often irrespective of their exact numerical values. However, in practice, these hyperparameters critically affect not only the convergence rate but also the computational efficiency of the solver. Techniques such as adaptive step-size selection or backtracking line search are commonly employed to accelerate convergence in terms of iterations [7, Ch. 3], [8]. Yet, these methods often increase per-iteration complexity, as they require evaluating multiple candidate updates, ultimately leading to increased overall latency, which is the very constraint that practical systems aim to minimize. Other methods include derivative-free optimization, where surrogate models and adaptive restart mechanisms are used to enhance robustness and efficiency under expensive or noisy function evaluations [9]. However, these solvers typically require multiple sequential objective evaluations and do not address the per-iteration computational cost, as they rely on full-precision model evaluations.
@@
Recent years have witnessed a growing interest in leveraging deep learning tools to enhance optimization algorithms in various applications [10, 11, 12], giving rise to the paradigm of learning to optimize [13]. Among the most structured and interpretable approaches within this paradigm is deep unfolding [14, 15], which systematically transforms iterative optimization algorithms into trainable neural network architectures. In deep unfolding, each iteration of the original solver is mapped to a layer in a feedforward network, with a common approach treating its algorithmic hyperparameters as learnable parameters [16]. A key advantage of this approach lies in its ability to optimize the optimizer itself under a fixed computational budget, allowing one to preserve the interpretability and domain-specific structure of the original method while tuning its behavior to match the statistics of the data [17]. This has been shown to enable reliable operation with a limited number of iterations of unfolded optimizers adopted in various domains, including wireless communications [18, 19], distributed optimization [20, 21], beamforming [22, 23, 24], integrated sensing and communications [25, 26], and outlier detection [27, 28], making the method highly suitable for latency-sensitive applications. Moreover, since the learning process is performed offline, no additional computational overhead is introduced at inference time, unlike adaptive schemes such as backtracking.
@@
In this paper, we propose a new class of learned optimization solvers based on deep unfolding, which compensates both for a limited number of iterations and for approximated computations within each iteration. Beyond conventional deep unfolding, which primarily parametrizes iterative solvers [14], this work establishes deep unfolding as a principled framework for complexity reduction. This approach enables the deliberate incorporation of low-complexity approximations within iterative updates and compensating for their induced distortions through learnable parameterization. Our key insight is that leveraging data allows designing unfolded architectures that effectively absorb the performance loss introduced by deliberately simplified operations. This enables designing solvers that are both fast, due to a fixed and small number of unfolded iterations, and lightweight, due to replacing computationally expensive steps with efficient surrogates. Unlike standard deep unfolding methods that focus on reducing iteration count via, e.g., learned hyperparameters, our framework introduces a second axis of approximation: replacing costly operations within each iteration with low-complexity surrogates. These approximations are compensated for by expanding the learnable parameter space, allowing the optimizer to adaptively absorb their effects. This dual approximation (in both iteration depth and per-iteration complexity) enhances our approach’s applicability in latency- and resource-constrained environments. In doing so, we uncover an additional core benefit of deep unfolding: its capacity to tolerate and learn around structural approximations, thereby reducing both the number and cost of iterations without sacrificing solution quality.
@@
  • A new paradigm for learned approximated optimization: We introduce a principled framework for constructing unfolded approximated iterative optimizers, wherein selected steps of a classical iterative solver are replaced with lower-complexity operations. These approximate steps are embedded in a fixed-depth unfolded network whose parameters are optimized using training data, allowing the learned model to best match the data distribution.
@@
  • Case study II — Robust principal component analysis: We further demonstrate the versatility of our approach by applying it to RPCA [30], a fundamental problem in signal processing. Using an unfolded version of the iterative thresholding algorithm with approximated gradient computations, we design a learned RPCA solver that maintains competitive recovery accuracy while dramatically reducing runtime.
@@ II-A Optimization Model
where $h_{{\bm{\theta}}^{\rm h}_{k}}:\mathcal{S}\times\mathcal{X}\mapsto\mathcal{S}$ denotes the update function applied at iteration $k$. This update is governed by solver-specific parameters ${\bm{\theta}}^{\rm h}_{k}$ (e.g., step-sizes or momentum factors), referred to as hyperparameters, which guide the update steps but do not alter the optimization objective itself. While often fixed to constant values, which are either hand-tuned or derived via theoretical convergence criteria, hyperparameters play a critical role in determining the convergence speed, runtime and overall solution quality of the optimizer.
@@ III-A Rationale
As noted in C1 and C2, a major bottleneck in applying iterative optimization solvers to latency-sensitive applications stems from the variable and potentially large number of iterations required for convergence. To support rapid decision-making, it is thus essential to design optimization algorithms that operate with a fixed and relatively small number of iterations $K$, chosen in advance. A key enabler for such fixed-depth operation is the framework of deep unfolding, which has been shown to preserve the interpretable structure of classical solvers while learning iteration-specific hyperparameters from data [17]. In this paradigm, the iterative mapping (3) is viewed as a $K$-layer neural network, with each layer corresponding to a single iteration whose parameters are tuned offline for the target domain. Interpretability here follows from the optimization-based structure, as each layer corresponds to an optimization step and produces an intermediate estimate of the decision variables. Consequently, the internal representations exchanged between layers correspond to progressively refined estimates of $\bm{s}$}. By casting the hyperparameters $\{{\bm{\theta}}^{\rm h}_{k}\}_{k=1}^{K}$ as the trainable parameters of a discriminative machine learning model [31], one fully retains the interpretable operation of the original solver while achieving reliable performance within a small number of steps.
However, limiting the number of iterations alone does not fully resolve the latency bottleneck. As discussed in C3, the complexity of each iteration may itself be prohibitive, particularly when involving costly gradient evaluations or matrix operations. To address this, we take a further step by replacing selected iterations with low-complexity approximated computations, designed to significantly reduce the per-iteration runtime. Of course, such approximations may degrade performance if left uncompensated. To maintain the reliability of the resulting optimization process, we extend the hyperparameter space of the unfolded solver in a way that does not increase computational burden (e.g., by using element-wise or structure-aware parameterizations), and we train the full approximated optimization procedure end-to-end using data. In this manner, we obtain an optimization model that is both interpretable and efficient, and can achieve strong performance with low latency and low computational overhead.
@@ III-B1 Increased Parameterization
This formulation increases the flexibility in learning iteration-specific dynamics, without increasing the number of floating-point operations, as the element-wise product is of the same computational order as scalar multiplication. This approach is reminiscent of variable metric algorithms with diagonal preconditioners. In addition, it also preserves the semantics of descent-based optimization under mild regularity assumptions, as stated in the following proposition.
@@
where the approximation in $(a)$ holds for sufficiently small step-sizes, and $(b)$ is obtained by substituting the update step from (5) and rearranging, with ${\rm diag}(\bm{\eta}_{k})$ being a diagonal matrix whose diagonal entries are $\bm{\eta}_{k}$. Since these entries are positive, then this diagonal matrix is positive-definite, and thus, $\mathcal{L}_{\rm o}({\bm{s}}^{(k+1)};{\bm{x}})\leq\mathcal{L}_{\rm o}({\bm{s}}^{(k)};{\bm{x}})$, i.e., (5) is a descent method. ∎
Proposition 1 demonstrates that, using our running example of gradient descent, increased abstractness through element-wise parameterization can be made to $(i)$ preserve computational complexity; $(ii)$ not compromise theoretical guarantees such as descent behavior; and $(iii)$ provide richer representational power to absorb approximations. It does not provide finite-iteration performance guarantees; rather, performance within a fixed number of iterations $K$ is achieved via the data-driven training procedure detailed in the sequel.
@@ III-B2 Constructing the Approximated Iterative Optimizer
Assume that $\mathcal{L}_{\rm o}({\bm{s}};{\bm{x}})$ is $\mu$-strongly convex and $L$-smooth with $L\geq\mu>0$. Consider the recursion in (7), and let the step-size vector $\bm{\eta}_{k}$ satisfy $0<\underline{\eta}_{k}\leq[\bm{\eta}_{k}]_{i}\leq\bar{\eta}_{k}<\frac{2}{L}$ for all $i$, and that for every $k\in\mathcal{K}^{\rm approx}$,
@@
while for $k\notin\mathcal{K}^{\rm approx}$ we set $\delta_{k}=0$. Then, for any minimizer $\bm{s}$$^{\star}$ of $\mathcal{L}_{\rm o}(\cdot;{\bm{x}})$, the objective after $K$ steps satisfies:
$\mathcal{L}_{\rm o}({\bm{s}}^{(K)};{\bm{x}})\!-\!\mathcal{L}_{\rm o}({\bm{s}}^{\star};{\bm{x}})\leq$ $(1\!-\!\mu c)^{K}\!\left(\mathcal{L}_{\rm o}({\bm{s}}^{(0)};{\bm{x}})\!-\!\mathcal{L}_{\rm o}({\bm{s}}^{\star};{\bm{x}})\right)$
$+\sum_{k=0}^{K-1}(1-\mu c)^{K-k-1}C\,\delta_{k}^{2},$ (9)
where $c\triangleq\inf_{k}\underline{\eta}_{k}\!\left(1-\frac{L}{2}\sup_{k}\bar{\eta}_{k}\right)$, $C\triangleq\frac{L}{2}+\frac{(1+L\sup_{k}\bar{\eta}_{k})^{2}}{2c}$.
@@
Proposition 2 characterizes how approximation errors accumulate in the gradient-descent example and how their impact is modulated by the step-size parameters. Specifically, the first term in (9) corresponds to the standard convergence rate of gradient descent [32, Thm. 3.4]. The second term in (9) accumulates errors that are based on the deviations from the true gradient only over the selected iterations $\mathcal{K}^{\rm approx}$. This result shows that the final error scales with the magnitude of the skipped gradients and the approximation error $\delta_{k}$. The proposition indicates the potential usefulness of deliberately inducing approximated computations (for complexity reduction) with increased parameterization, since as long as the skipped gradients (balanced by the increased hyperparameters $\bm{\eta}_{k}$) are small, the error remains bounded and the method remains effective. Proposition 2 provides theoretical insight into the interplay between approximation and parameterization, rather than guaranteeing bounded error for arbitrary approximation levels; using deep unfolding techniques, the hyperparameters are learned from data to compensate for approximation-induced distortions, as detailed in the following subsection.
@@ III-B3 Training via Empirical Risk Minimization
To learn the parameters $\Theta\triangleq\{\Theta_{k}\}_{k=1}^{K}$ of the approximated optimizer, we assume access to a representative dataset $\mathcal{D}$. The training procedure depends on the nature of the available supervision.
@@
where $\mathcal{L}_{\rm task}(\cdot,\cdot)$ is a supervised loss function chosen according to the target application. Supervised learning is most useful in settings where the optimization objective $\mathcal{L}_{\rm o}$ is a mathematically formulated proxy of the actual system task.
In both cases, the unfolded approximated optimizer is trained offline using standard gradient-based methods to minimize the empirical risk over $\Theta$. This allows the model to compensate for the approximations in both iteration count and complexity, yielding an efficient and reliable learned decision rule. The ability of empirical risk minimization to compensate for the performance loss introduced by computational approximations stems from two main factors: $(i)$ the increased parameterization compared to the original iterative solver, which provides additional degrees of freedom that enable the learned optimizer to minimize the influence of approximation error terms in the error-bound expansion (as shown in Proposition 2 for gradient descent); and $(ii)$ the parameters are learned end-to-end via empirical risk minimization, where the loss evaluates the quality of the final output of the $K$-layer architecture, This allows the training process to identify parameter configurations that learn the implicit bias of the data distribution, effectively ’weighting’ the update steps to correct distortions in the intermediate update trajectory caused by the approximated computations. The formulation of training an approximated unfolded optimizer via conventional deep learning methods, i.e., mini-batch sgd (sgd), is stated in Algorithm 1.
@@ III-C Discussion
The proposed methodology introduces a novel and flexible approach to accelerate iterative optimization by combining deep unfolding with approximated computations. While conventional uses of deep unfolding aim to reduce the number of iterations by learning solver hyperparameters from data, our framework extends this idea by deliberately introducing low-complexity approximations within the iterations themselves. These approximations are typically regarded as limitations that degrade performance; however, we show that when integrated into an unfolded architecture and paired with appropriate parameterization, the resulting model can be trained to absorb and compensate for the induced mismatches. In doing so, our methodology directly addresses all three challenges C1-C3. This dual approximation on both iteration depth and per-iteration complexity is made tractable and effective through the trainability offered by deep unfolding. The increased abstractness in adding hyperparameters is designed not to come at the cost of increased complexity. Moreover, different abstractions, e.g., the usage of a per-parameter step-size rather than scalar ones as we do in our case studies, in fact preserve the property of descent methods as in (2[33].
Like most learning-based methods, unfolded optimizers may require retraining when the system configuration changes, particularly when their parameters depend on the data size (as is the case with extended hyperparameters). However, due to their structured origin in iterative optimization algorithms, unfolded architectures often exhibit stronger scalability properties than generic deep neural networks. In particular, as the trainable parameters correspond to solver hyperparameters, these parameters can often be extended across problem dimensions through simple scaling or replication strategies [20] to enable zero-shot generalization to larger problem instances. Alternatively, the modular structure of unfolded optimizers can be combined with hypernetworks that generate hyperparameters conditioned on the system configuration as proposed in [34], enabling elastic architectures capable of adapting to varying problem dimensions without retraining. These extensions are left for future study.
The proposed framework is intended as a design methodology rather than an off-the-shelf ‘plug-and-play’ algorithm. Because the sources of computational complexity vary significantly across different iterative solvers applications, the choice of which operations to approximate and how to approximate them must be tailored on a case-by-case basis. Accordingly, many of the practical considerations and benefits of our framework are revealed only when applied to specific use cases. In the following sections, we illustrate how our general methodology can be instantiated in two representative signal processing tasks. These case studies demonstrate how approximations can be selected, quantify their computational impact, and empirically validate the ability of the learned optimizer to preserve or even improve performance.